reduce the given loading system to a force couple system at point a. then determine the distance x to the right of a point a at which the resultant of the three forces acts
1
Expert's answer
2021-10-15T10:22:17-0400

"F_{eq} = -800+120-720 = -321 \;N \\

M_A = 1200 \times 450 -720 \times 700 -800 \times 200 = -12400 \;N \cdot mm"

Using Varignon’s theorem

"F_{eq} \times x = M_A \\

x = \frac{124000}{320} = 387.5 \;mm"

Answer: 387.5 mm